### v2

# Arrows of Time for Large Language Models

Abstract
We study the probabilistic modeling performed by Autoregressive Large Language Models through the angle of time directionality. We empirically find a time asymmetry exhibited by such models in their ability to model natural language: a difference in the average log-perplexity when trying to predict the next token versus when trying to predict the previous one. This difference is at the same time subtle and very consistent across various modalities (language, model size, training time, …). Theoretically, this is surprising: from an information-theoretic point of view, there should be no such difference. We provide a theoretical framework to explain how such an asymmetry can appear from sparsity and computational complexity considerations, and outline a number of perspectives opened by our results.
Machine Learning, ICML
1 Introduction
Generative Models have revolutionized modern AI, yielding a wide array of applications. Modern works have shown that such models can perform spectacularly (and somewhat mysteriously) well on various kinds of data. Text is perhaps the domain where progress has been the most drastic: in a few years, Large Language Models (LLMs) have gone from generating barely correct sentences to producing consistent stories, code, and performing countless new tasks; key milestones include the Transformer architecture (Vaswani et al., 2017), BERT (Devlin et al., 2019), and GPTs (Radford et al., 2018, 2019; Brown et al., 2020; OpenAI, 2023).
At the heart of these developments are probabilistic models trained in an unsupervised manner on vast amounts of data, for prediction or recovery tasks: this yields an estimation of the probability measure underlying the data. These probabilistic models appear to gain surprising abilities, such as reasoning, as their sizes increase (see (Wei et al., 2022; Schaeffer et al., 2023) among others).
In this work, we investigate the interplay between the probabilistic structure of autoregressive LLMs and the data they are trained on. More precisely, we investigate how time directionality influences their ability to model natural and synthetic languages.
1.1 Autoregressive LLMs
Famously, the pre-training of LLMs such as the GPTs consists in ‘learning to predict the next token’ knowing previous ones, in sequences extracted from large text corpora, using the natural time ordering of the data they are being trained on. A vocabulary $\mathcal{V}$ of $V$ tokens is chosen; the dataset is then tokenized into a sequence of tokens in $\mathcal{V}$; at each step, the model reads a sequence of tokens and outputs a probability distribution on $\mathcal{V}$ predicting the next token.
Typically, as a probabilistic model, an autoregressive model will estimate the probability that $n$ random consecutive tokens $\left(X_{1},\cdots,X_{n}\right)$ are equal to $\left(x_{1},\cdots,x_{n}\right)\in\mathcal{V}^{n}$ by taking the product of the (estimations of) the probabilities
$\mathbb{P}$$\left\{X_{1}=x_{1}\right\}$
$\mathbb{P}$$\left\{X_{2}=x_{2}|X_{1}=x_{1}\right\}$
$\vdots$ (1)
$\mathbb{P}$$\left\{X_{n}=x_{n}|X_{1}=x_{1},\cdots,X_{n-1}=x_{n-1}\right\},$
yielding an estimated probability measure $\mathbb{P}_{n}^{\rightarrow}$ on $\mathcal{V}^{n}$.
Autoregressive LLMs (GPTs, GRUs, LSTMs, …) thus factorize (their estimates of) the joint probabilities in terms of conditional probabilities for each token knowing past ones. This brings a number of advantages: first, leverage the fact that for each token sequence $x_{1},\ldots,x_{n}$, each token $x_{k}$ is used in to predict each token $x_{\ell}$ with $\ell>k$. In particular, GPT (compared to the earlier BERT) includes causality-aware attention, allowing for a parallelization of the training process: a sequence $x_{1},\ldots,x_{n}$ generates $n-1$ fully parallelizable tasks (predict $x_{k}$ from $\left(x_{1},\ldots,x_{k-1}\right)$ for $2\leq k\leq n$). Also, this representation enables a natural sampling from $\mathbb{P}_{n}^{\rightarrow}$ (token by token), as well as data compression: the factorization decomposes these processes into many smaller substeps (Graves et al., 2023).
Autoregressive LLMs such as GPTs have enabled a massive scaling up of the number of parameters and dataset sizes, yielding numerous fascinating phenomena, e.g. scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022) and emergent behavior, e.g. abilities at arithmetic operations (Shen et al., 2023), circuit computing tasks (d’Ascoli et al., 2023), or high-level linguistic proficiency.
1.2 Arrow of Time and Language Models
While decomposing measures into a sequence of conditional probabilities is natural, it is not a priori clear why following the time direction of language to do so is optimal (except for downstream tasks, e.g. making a chatbot): what is the best order when predicting the token probabilities? A natural idea to investigate this question is simply to reverse the Arrow of Time: to estimate probabilities backward. This amounts to training models on time-flipped datasets: we train the same models on the same data slices for next-token predictions, but for each data slice $\left(x_{1},\cdots,x_{n}\right)$ we feed the model with $\left(x_{n},x_{n-1},\cdots,x_{1}\right)$ instead.
As a result, instead of (1), we take the product of the estimations of
$\mathbb{P}$$\left\{X_{n}=x_{n}\right\}$
$\mathbb{P}$$\left\{X_{n-1}=x_{n-1}|X_{n}=x_{n}\right\}$
$\vdots$ (2)
$\mathbb{P}$$\left\{X_{1}=x_{1}|X_{n}=x_{n},\cdots,X_{2}=x_{2}\right\}.$
This yields an estimated probability measure $\mathbb{P}_{n}^{\leftarrow}$ on $\mathcal{V}^{n}$.
In this paper, we will speak of forward/backward (FW/BW) model to refer to the same (architectural) model trained with the same hyperparameters (learning rate, batch size, training time, …) but fed with (batches of) $\left(x_{1},\ldots,x_{n}\right)$ and $\left(x_{n},\ldots,x_{1}\right)$ from the same dataset respectively. In other words, both models are the same, except that the FW model is trained to predict the next token, while the BW one is trained to predict the previous token.
Problem 1.
For a measure $\mathbb{P}$ and a given model, how do the forward and backward measures $\mathbb{P}_{n}^{\rightarrow}$ and $\mathbb{P}_{n}^{\leftarrow}$ differ from one another?
For certain $\mathbb{P}$s, we will see universal asymmetries: for any given architecture and hyperparameters, a substantial difference between the way $\mathbb{P}_{n}^{\rightarrow}$ and $\mathbb{P}_{n}^{\leftarrow}$ approximate $\mathbb{P}$ arises.
1.3 Cross-Entropy Loss and Perplexity
LLMs are trained as follows: sample sequences of $n$ consecutive tokens $\left(x_{1},\ldots,x_{n}\right)$ from the dataset; then, for $i=1,\ldots,n$, get a prediction $p_{i}:\mathcal{V}\to\left[0,1\right]$ for $X_{i}$ (given previous tokens for the FW model/next tokens for the BW one), compute the loss $\sum_{i=1}^{n}\ell\left(p_{i},x_{i}\right)$ on the observed tokens $x_{i}$ for a loss function $\ell$, perform a gradient step to optimize $\ell$, and start again.
In the training of most LLMs, the prime choice for $\ell$ is the cross-entropy loss, defined by $\ell_{\mathcal{C}}\left(\mathbf{p}_{k},x_{k}\right)=-\ln\mathbf{p}_{k}\left(x_{k}\right)$: the negative log of the predicted probability of the observed token. It is a proper scoring rule (Savage, 1971; Gneiting & Raftery, 2007), uniquely identified by certain modularity properties (Hanson, 2012); in expectation, it gives the number of nats ($\ln{2}$ times the number of bits) needed to compress $\left(x_{1},\ldots,x_{n}\right)$ when using a coding scheme based on the model’s estimated probabilities. Finally, and crucially for us, we have the following:
Remark 2.
For $i=1,\ldots,n$, let $\left(\mathbf{p}_{i}^{\rightarrow}\right)_{i}$ and $\left(\mathbf{p}_{i}^{\leftarrow}\right)_{i}$ denote the predictions of the FW and BW models respectively. Setting $\ell_{i}^{\leftrightarrows}:=\ell_{\mathcal{C}}\left(\mathbf{p}_{i}^{\leftrightarrows},x_{i}\right)$, we have
$\sum_{i=1}^{n}\ell_{i}^{\rightleftarrows}=-\ln\mathbb{P}_{n}^{\leftrightarrows}\left\{X_{1}=x_{1},\cdots,X_{n}=x_{n}\right\}.$
In particular, if the FW and BW measures coincide, the cross-entropy losses are identical.
Remark 3.
If $\left(x_{1},\ldots,x_{n}\right)$ is sampled from $\mathbb{P}_{n}$, denoting by $\mathcal{L}_{n}^{\leftrightarrows}$ the expectations of $\sum_{i=1}^{n}\ell_{i}^{\leftrightarrows}$ (estimated by the test loss of the models during training), we have
$\mathcal{L}_{n}^{\leftrightarrows}=\mathrm{D}_{\mathrm{KL}}\left(\mathbb{P}_{n}\big{|}\big{|}\mathbb{P}_{n}^{\leftrightarrows}\right)+H\left(\mathbb{P}_{n}\right),$
where $H$ denotes the entropy and $\mathrm{D}_{\mathrm{KL}}$ the Kullback-Leibler divergence.
Remark 4.
It is worth noting that, in spite of its apparent triviality, Remark 2 crucially depends on the choice $\ell$ as $\ell_{\mathcal{C}}$. Moreover, even if $\mathbb{P}_{n}^{\rightarrow}=\mathbb{P}_{n}^{\leftarrow}$, we will generally have $\ell_{i}^{\rightarrow}\neq\ell_{i}^{\leftarrow}$ for $1\leq i\leq n$ (though $\sum_{i}\ell_{i}^{\rightarrow}=\sum_{i}\ell_{i}^{\leftarrow}$). When $\mathbb{P}_{n}^{\rightarrow}=\mathbb{P}_{n}^{\leftarrow}=\mathbb{P}_{n}$ we have that $\ell_{i}^{\rightarrow}$ and $\ell_{i}^{\leftarrow}$ yield two (typically different) decompositions of the log-likelihood of $\left(x_{1},\ldots,x_{n}\right)$. For instance, take as a dataset the 81 expressions $A\times B=CD$ for $1\leq A,B\leq 9$ (setting $C=0$ when needed). The FW log-perplexity is concentrated on $A$ and $B$, each contributing $\ln{9}\approx 2.2$ nats: $\ell_{A}^{\rightarrow}=\ell_{B}^{\rightarrow}\approx 2.2$, $\ell_{C}^{\rightarrow}=\ell_{D}^{\rightarrow}=0$. The BW log-perplexity is distributed differently: for instance, for $3\times 4=12$, $(\ell_{A}^{\leftarrow},\ell_{B}^{\leftarrow},\ell_{C}^{\leftarrow},\ell_{D}^{\leftarrow})\approx(0,1.39,1.1,1.91)$.
Remark 5.
Remark 3 suggests that if $\mathbb{P}_{n}^{\rightarrow}$ and $\mathbb{P}_{n}^{\leftarrow}$ coincide (e.g. if both models have learned the true measure $\mathbb{P}$, memorized the training set, or more generally have learned $\mathbb{P}$ ‘equally well’), their associated average losses should be equal. If we take a very small dataset or context length, we can expect to have $\mathcal{L}_{n}^{\rightarrow}\approx\mathcal{L}_{n}^{\leftarrow}$. If we are to train a FW and a BW model with our setting, any substantial difference in their cross-entropy losses will necessarily reflect an asymmetry of $\mathbb{P}$ (w.r.t. its learnability by the models).
As it will turn out, for many types of data (i.e. $\mathbb{P}$s), a consistent difference between FW and BW log-perplexities arises across a wide range of models and hyperparameters.
1.4 Setup and Plan
In Section 1.3, we showed that a difference between FW and BW losses reflects a difference between the measures $\mathbb{P}_{n}^{\rightarrow}$ and $\mathbb{P}_{n}^{\leftarrow}$ learned by the FW/BW models, all else being equal: same dataset, same model (architecture and hyperparameters). In such a case, we say that $\mathbb{P}$ (or the corresponding dataset) exhibits an Arrow of Time (AoT) with respect to the model and context length $n$. We speak of a FW AoT if the average FW log-perplexity is below the BW one (i.e. if the FW model outperforms the BW one).
This paper aims to investigate the following questions:
  • Is there an AoT in large natural language datasets? Does it depend on the language? Does it depend on the context length $n$?
  • Can we formulate a theoretical framework explaining the presence of an AoT? Can we construct simple synthetic datasets exhibiting an AoT? Can the presence of an AoT be explained mathematically from first principles? What should we expect as the model sizes tend to infinity?
The paper is organized as follows:
  • In Section 2, we investigate the existence of an AoT, starting from a basic setup and expanding across modalities: languages, architectures, hyperparameters, context lengths.
  • In Section 3, we investigate the theoretical origins of AoTs, starting with a simple synthetic dataset exhibiting one; in this case, the difference between $\mathcal{L}_{n}^{\rightarrow}$ and $\mathcal{L}_{n}^{\leftarrow}$ can be shown to be related to the hardness of the factoring problem (Section 3.1). We then introduce a more general class of synthetic datasets which we call ‘linear languages’ (Section 3.2), providing us with a fairly wide class of datasets with a mathematically justified AoT.
  • Finally, in Section 4, we summarize our results and outline a number of possible future research directions.
1.5 Relation to Previous Works
The idea of training language models ‘backward’, on reverse sequences, has occurred in a number of contexts. A notable example is (Sutskever et al., 2014), focusing on machine translation using LSTMs, finding that reversing the source sentence (i.e. the source LSTM learning backward) greatly improves performance. Other well-known examples of related techniques include ULMFiT (Howard & Ruder, 2018) and ELMO (Peters et al., 2018).
Attempts at combining FW/BW models include (Mou et al., 2016; Liu et al., 2016; Zhang et al., 2018; Serdyuk et al., 2018; Mangal et al., 2019) (using RNNs); or recently (Nguyen et al., 2023), a ‘Meet in the Middle’ approach which shows how pre-training using FW/BW Transformer models enhance FW-only autoregressive generation; as well as in (Shen et al., 2023), applying the idea of reversing data to improve LLM performance (see also Section 3.1). In these approaches, the FW/BW models are treated as one model, yielding one combined loss. They compare various models on a task (see section 5 of (Nguyen et al., 2023) for a comprehensive review), rather than study potential discrepancies in FW/BW learning, as in most papers combining FW and BW generation.
A recent paper on adversarial attacks on LLMs (Pfau et al., 2023) compares FW/BW LLMs (with a BW GPT, 160M parameters), and an older study (Duchateau et al., 2002), uses BW N-gram language models for continuous speech recognition (comparing FW/BW trigrams models). In those, the authors report that BW models perform respectively equally or slightly better than FW models, in contradiction with our main findings. This does not affect our confidence in our results, given the magnitude of our experiments and the level of care involved in their setup.
2 Empirical Results on Natural Language
In this section, we explore the existence of an AoT for LLMs in natural language datasets. We first reveal the presence of an AoT in a basic setup (GPT2 models on English and French with context window of length 256, see Section 2.2.1 below). We then decline our explorations over more than 50 model modalities (GPT/GRU/LSTM architectures, 6 GPT sizes, context window lengths of 16/32/64/128/256/512 and 8 languages), and rule out possible tokenization artifacts. We observe a consistent FW AoT in these setups, including a number of takeaways concerning its magnitude.
2.1 Setup
For the identification of an AoT in a dataset, we make sure that both the FW and BW models are trained with the exact same specifications: the only reason for a difference between the models’ performances is hence the token prediction order. In all experiments, the models are trained from scratch, using He initialization.
2.1.1 Dataset and Tokenization
[…]
2.1.2 Models, Hyperparameters and Training
While some experiments involve other autoregressive models (GRU, LSTM), for most training jobs we use the decoder-only Transformer (GPT) (Radford et al., 2018); our implementation (with code in the supplementary material) is derived from minGPT (Karpathy, 2020). All GPT experiments use learned positional embeddings and a dropout rate of $0.1$. Other hyperparameters may depend on the experiment, see the Appendix.
For all models, we use the AdamW optimizer (Loshchilov & Hutter, 2019) with base learning rate of $10^{-4}$ and a learning rate schedule with a warmup, followed by cosine annealing with warm restarts (Loshchilov & Hutter, 2017). These hyperparameters are mostly kept constant across different experiments, although the period of the warm restarts might be tweaked to synchronize the end of training with the end of a cycle, see the Appendix.
2.2 Results
In this section, we present results of various experiments confirming the presence of a FW AoT in English and French datasets, and provide compelling evidence for its universal existence in natural languages by considering six other languages (five distinct families in total).
2.2.1 Arrow of Time in English and French
We begin in this section by analyzing the difference in FW vs BW training dynamics for a Transformer of size GPT2-Medium ($\sim 405M$ parameters, context window length of $256$) on the CC100 datasets for English and French. We train the FW and BW models for the equivalent of 1 epoch of the French dataset ($\sim 27B$ tokens), avoiding memorization.
Figure 1: English vs French validation losses (French training losses in the zoom-in, early loss values cropped for readability).
As is seen in the zoom-in of Fig 1., after an initial short transition period, the BW model loss separates from its FW counterpart and settles slightly above it, and then follows an almost parallel trajectory. This consistent difference throughout training (even persisting through warm restarts) points to the existence of an AoT both in English and in French: at the end of training, we see the following losses for English: FW: $2.88$, BW: $2.902$ a difference of $+0.76\%$; and for French: FW: $2.788$, BW $2.862$, a difference of $+2.65\%$. Interestingly, the magnitude of this effect is different for English and French.
As will be discussed in the next subsections, the findings are quite universal: they can be consistently expanded to various settings, across models, languages, and context lengths.
2.2.2 Context Window Size
In this section, we examine the influence of long-range correlations on the AoT, by studying its relationship with the context length. Intuitively, for a very small context length, we should see virtually no AoT; with very few tokens, models approach the optimal solutions similarly, as they have fewer degrees of freedom. For instance, in the extreme case of a context of length 2, models are only tasked with learning a two-variable function $\mathcal{V}^{2}\to\left[0,1\right]$, i.e. to learn the frequencies of 2-grams, which should be (equally) easy in both directions. It is likely that an AoT emerges for larger context lengths (and for reasonably large models).
We test the dependence on the context window by training the same GPT-Medium model, but with context lengths spanning from $16$ to $512$ tokens, both on English and French. As can be seen in Fig. 2, the magnitude of the AoT in both English and French increases with the context size, suggesting the importance of long-range dependences.
Figure 2: BW/FW losses percentage difference for different context lengths
2.2.3 Model Size
In this section, we investigate the effect of model size for GPT models (other models are discussed in the next subsection). As in Section 2.2.2 above, it is natural to expect small models to struggle to exhibit an AoT that would depend on sophisticated, long-range dependences. To test this, we train GPT models of different sizes, from $5M$ to $405M$ parameters, all with a context length of $256$. Interestingly the AoT is much smaller at very small model sizes, reinforcing the idea that long-range dependences are key; as the model size keeps growing beyond that, the difference tends to grow. Note also that larger BW models typically outperform smaller FW ones.
Table 1: Final FW losses and relative BW differences.
Nano Micro Mini Small GPT1 Med
Size 4.92M 13.7M 22.0M 55.6M 162M 405M
Fr-FW 4.525 3.964 3.683 3.293 2.979 2.788
Fr-BW +0.15% +0.63% +1.49% +1.64% +2.07% +2.65%
En-FW 4.599 4.064 3.799 3.416 3.081 2.880
En-BW -0.33% +0.1% +0.11% +0.26% +0.49% +0.76%
2.2.4 Other Models
While most results in this paper are focused on GPT models (the current state of the art for language modeling), the question of the AoT can naturally be asked for other autoregressive models. We investigate this for GRUs and LSTMs (three sizes each), again with a context length of 256.
Once more, for sufficiently large models, we observe a consistent AoT throughout modalities, confirming that the observed AoT goes beyond Transformer models; rather, it appears to be intrinsic to the dataset. It is interesting for instance that for the English dataset, the smaller BW model performs slightly better than the FW one. This however convincingly disappears for larger context sizes and models.
Table 2: Final FW GRU/LSTM losses and relative BW differences.
GRU S GRU M GRU L LSTM S LSTM M LSTM L
Size 4.92M 13.7M 22.0M 55.6M 162M 405M
Fr-FW 3.905 3.692 3.363 3.901 3.566 3.314
Fr-BW +0.26% +0.3% +0.62% +0.1% +0.45% +0.66%
En-FW 4.030 3.712 3.483 4.015 3.653 3.418
En-BW -0.07% +0.22% +0.34% -0.27% +0.11% +0.15%
2.2.5 Other Languages
The above experiments confirm the existence of an AoT for English and French across various modalities. An exciting question that naturally arises is whether this might be a universal property of natural languages. To begin to explore this question, we train models of two sizes (GPT2-Medium and GPT2-XL) on six more languages.
Table 3: Final losses for different languages, Medium/XL models. Format: [Final FW loss]/[BW relative difference].
German Turkish Finnish Vietnamese
Med 3.148/+1.46% 3.292/+0.94% 3.359/+1.07% 2.396/+1.67%
XL 2.892/+1.59% 3.084/+1.17% 2.975/+1.85% 2.099/+2.81%
Greek Indonesian French English
Med 2.794/+2.4% 3.079/+2.18% 2.834/+2.4% 2.926/+0.61%
XL 2.494/+3.05% 2.741/+3.17% 2.586/+2.51% 2.683/+0.63%
From Table 3, we can see that in all the cases we tested, a FW AoT emerges, although its magnitude appears to vary from language to language, suggesting some universality of this phenomenon across human languages.
2.2.6 Possible Tokenization Artifacts
While the training procedures are perfectly symmetric with respect to the two directions, it is important to rule out any other possible sources of asymmetry. One possible source could in principle be the tokenization; indeed, the BPE tokenizer is trained in the FW direction. To rule out this possibility, we inverted (at the character level) one dataset (Greek), and re-trained a BPE tokenizer on the result. We trained a GPT2-Medium on it, and confirmed that the direction of the BPE tokenization has no effect on the training dynamics: in this case, the FW (respectively BW) model performs exactly the same way as the BW (respectively FW) model on the original tokenization, thereby showing exactly the same AoT. See Appendix A for details.
2.3 Key Takeaways
The above experiments suggest the universality of the phenomenon of AoT across languages, models, and hyperparameters. More specifically:
  • A very consistent AoT emerges for large enough models, trained for long enough, and with a large enough context window; in the other cases, the effects are less clear.
  • An important finding is that the magnitude of the AoT increases with the context length: this suggests the importance of long-range correlations; relatedly, the model’s size can influence its ability to use the information of its whole context window.
  • While most of our training is done with GPT models, we observe the same type of results for GRUs and LSTMs, suggesting that AoTs are intrinsic to datasets.
  • An interesting phenomenon is that the magnitude of the AoT greatly depends on the language, even if its presence and direction are universal. Explaining this convincingly remains a fascinating challenge.
In Section 3, we introduce a framework to reveal the emergence of the AoT in synthetic datasets and propose mechanisms to explain how this can apply to natural languages. In Section 4, we discuss how these somehow surprising results open the door to many possible investigations.
3 Computability and Irreversibility
As discussed in 1.3 above, from an information-theoretic point of view (abstracting away computability), there should be no difference between FW/BW models. However, as shown in 2.2, we see a consistent AoT for various types of architectures across multiple modalities, which increases with larger context windows. As a result, any plausible explanation must explain why certain probabilities are harder either to be (1) represented or (2) learned with BW models than with FW ones. Naturally (1) is stronger than (2): models cannot learn what they cannot represent (e.g. if there exists no set of model weights that solve the problem). In this section, we provide simple mathematical models of data illustrating how both mechanisms can arise and naturally contribute to the AoT. We start with a simple mathematical model using prime number multiplications, illustrating how the computational hardness of reversing certain information-preserving operations generates an AoT. We then construct a more general class of data models based on binary operations, allowing one to reveal an AoT based on sparsity and complexity theory ideas.
3.1 Number Factoring and Arrow of Time
Perhaps the most classical example of information-preserving, yet hard to invert, computation is number factoring: given two large primes $p,q$ with $p<q$, it is relatively easy to compute $n=pq$; while $n$ contains the same information as $p$ and $q$, recovering them from their product is (believed to be) very hard. This problem is the basis of much of asymmetric cryptography. In this section, we study how FW/BW models perform when trained on a dataset based on this idea; we study the theoretical entropy distribution when reading the data FW and BW and compare this to the experimental values for FW/BW GPTs.
3.1.1 Synthetic dataset
For fixed $k\geq 1$, consider the synthetic language of strings of the form $p\times q\leftrightarrow\mathrm{rev}\left(pq\right)$, with $p<q$ primes, $p,q<10^{k}$ and $\mathrm{rev}\left(pq\right)$ being the product of $p$ times $q$, written in reverse order (see (Shen et al., 2023; Lee et al., 2023)). The numbers $p$ and $q$ are padded to be of $k$ digits exactly and the $\mathrm{rev}\left(pq\right)$ is padded to be of $2k$ digits exactly (e.g. for $k=4$: $0019\times 0023\leftrightarrow 73400000$). The symbols $\times$ and $\leftrightarrow$ are written as multiple tokens (3 tokens for $\times$, and 7 tokens for $\leftrightarrow$), to facilitate the learning of non-trivial operations by GPTs. For a fixed $k$, $\mathbb{P}$ is thus supported on sequences of $104k+$ 10 tokens; in our experiments, we set $k=5$ and take a dataset containing $10^{8}$ such random ordered pairs. Intuitively, computing the RHS of the symbol $\leftrightarrow$ given the LHS should be relatively easy (or at least learnable), while computing the LHS from the RHS should be much harder (at least finding the factor $q$; given $\mathrm{rev}\left(pq\right)$ and $q$, $p$ should be relatively easy to find).
3.1.2 Nats of entropy
In order to better understand the experimental results, we compute the aggregate entropy (in nats) on $p$, $q$ and $\mathrm{rev}\left(pq\right)$ when reading the strings $p\times q\leftrightarrow\mathrm{rev}\left(pq\right)$ FW and BW (we do not compute the entropy on each token individually, and the entropy on the symbols $\times$ and $\leftrightarrow$ is 00). For instance, for $k=5$, there are $\ln\left(\pi\left(10^{5}\right)\right)=9.17$ (with $\pi\left(x\right)=\#{p:p\leq x}$) nats of entropy over the possible prime numbers $<10^{5}$, which drops to $8.98$ nats of entropy on $p$ (because of the ordering), and (averaging over $p$) $8.67$ nats of entropy on $q$; this results in $17.64$ nats of entropy for the pair $\left(p,q\right)$, which is roughly $2\times 9.17-\ln\left(2\right)$ (we subtract the bit of information due to the ordering); since $\mathrm{rev}\left(pq\right)$ is determined by $p$ and $q$, its conditional entropy is naturally zero. Reading the string backward, the $17.64$ nats of entropy are concentrated on $\mathrm{rev}\left(pq\right)$; the rest is fully determined, and thus has zero entropy.
3.1.3 Experimental Results
Training a model with a GPT2-Medium on the $p\times q$ dataset yields the log-perplexities recorded in Table 4. The FW model is able to reach the information-theoretical limits on $p$ and $q$; the conditional cross-entropy loss on $\mathrm{rev}\left(pq\right)$ is low but non-zero, indicating that the model (imperfectly) learns to multiply the prime numbers. In contrast, the results for the BW model show a far-from-optimal perplexity on $\mathrm{rev}\left(pq\right)$, which points to the difficulty for the model to recognize the products of two primes; knowing $\mathrm{rev}\left(pq\right)$, almost no information on the prime factor $q$ is extracted: only $8.98-8.41=0.57$ nats, i.e. less than one bit. The ‘division’ is much more learnable, with all but $0.02$ nats of information learned. All in all, the total FW log-perplexity is $22.2$ nats, while the BW one reaches $30.2$ nats.
Table 4: Final perplexities for the prime numbers dataset
$p$ $q$ $rev(pq)$
FW 8.98 8.67 4.55
BW 0.02 8.41 21.56
3.1.4 Discussion
The above setup shows a significant AoT for the $p\times q\leftrightarrow\mathrm{rev}\left(pq\right)$ dataset. This discrepancy can be largely attributed to the asymmetry between the difficulty of factoring versus multiplication: compared to the information-theoretical limit, about $4.55$ nats are lost for the multiplication, while $8.43$ nats are lost for the factorization. We also see that the different structures of the LHS and RHS (which have the same information-theoretic content as they determine each other) also present a significant difference w.r.t. the models’ abilities: while the FW model reaches essentially optimal perplexity for the LHS (i.e. it recognizes primes $<10^{k}$), the BW model is very far from optimal on the RHS (i.e. to recognize products of primes pairs $p,q<10^{k}$ proves to be much more difficult). While part of the asymmetry is attributable to the models’ specifics, as long as the dataset size is kept so high that all pairs $\left(p,q\right)$ cannot be memorized, a significant AoT can be expected: as the model size (and training time) grows, multiplication will eventually be learned (long before the dataset can be memorized, (Shen et al., 2023)), while extracting substantial information from $\mathrm{rev}\left(pq\right)$ about $q$ should remain very hard. The above data model displays an AoT of types (1) and (2) (see 3 above): certain features turn out to be harder to learn for the BW models, while others simply are likely not even representable by LLMs (of reasonable sizes), as the alternative would yield an efficient factorization algorithm.
Remark 6.
Note also that the above also illustrates the importance of long-range dependences for the AoT: if the context length is kept e.g. significantly below $k$, the FW model will have trouble saying anything about $\mathrm{rev}\left(pq\right)$, as $p$ will already be forgotten when reaching the RHS, shrinking its advantage over the BW model.
Remark 7.
In the above setting, we are rooted in the computational difficulty of the inversion of a bijective function; note that the core of the argument is the computational difficulty, rather than bijectivity. Abstracting computability issues, there is still no difference between FW and BW perplexities for optimal predictors, even if a mapping is not injective, or if it is not well-defined as mapping (e.g. if some random noise is added to it). The invertibility merely helps us get a simple computation of the theoretical entropies, and to pinpoint where each model performs suboptimally; it is however not directly related to the presence of an AoT.
3.2 Binary Operations
The model of Section 3.1 shows how an AoT can appear in a dataset: in that example, based on computational complexity ideas, we could handcraft a synthetic dataset that is both practically and theoretically harder for a BW model than a FW one. This still leaves the question of why an AoT would arise in a dataset such as natural language, as in Section 2, and why in one direction rather than another, i.e. why FW models would consistently outperform BW ones. In this section, we introduce synthetic datasets based on operations on the space of $m$-bit registers identified with $\mathbb{F}_{2}^{m}$ ($\mathbb{F}_{2}$ denotes the field of integers mod $2$). We focus on languages based on $\mathbb{F}_{2}$-linear circuits and relate their learnability to their sparsity and use this to provide an explanation for a difference between FW and BW learnability; we then provide a setup motivating the specific FW direction observed on the AoT in natural languages; we then provide experiments validating our framework; we conclude by discussing extensions to the nonlinear case.
3.2.1 Linear Sparse Circuit Dynamics
Consider the class of measures $\mathbb{P}_{n}$ on a linear language formed by sequences of $n=2m+1$ tokens, of the form $x\leftrightarrow y$ where $x,y$ are random uniform on $\mathbb{F}_{2}^{m}$, but related by a bijective linear map; $\leftrightarrow$ is counted as a token. For each $\mathbb{P}_{n}$, we can write $y=f_{\mathbb{P}_{n}}^{\rightarrow}\left(x\right)$ and $x=f_{\mathbb{P}_{n}}^{\leftarrow}\left(y\right)$ for $f^{\leftrightarrows}:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2}^{m}$. We define the sparsity of a linear map $f:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2}^{m}$ as the proportion of zero entries of its matrix. We will (informally) say that a matrix is sparse if this proportion is relatively high, i.e. close to $1$. Intuitively, the sparsity of $f^{\rightarrow}$ (resp. of $f^{\leftarrow}$) is related to how easy it is to learn $\mathbb{P}_{n}$ (based on random data samples) for a FW model (resp. for a BW model). For GPT predictors, this is studied numerically in Section 3.2.3 below. For a linear language $\mathbb{P}_{n}$, an AoT will thus emerge if the sparsities of $f^{\rightarrow}$ and $f^{\leftarrow}$ are significantly different. It is common knowledge that the inverse of a sparse matrix is generally less sparse (see e.g. (Duff et al., 2017), Section 15.6), This is the basis for the following claim (verified numerically in Appendix B.2.2):
Claim 8.
If $A$ is a sparse random $m\times m$ matrix in $\mathbb{F}_{2}$ conditioned to be invertible, the matrix $A^{-1}$ has typically lower sparsity. Similarly, if we perturb a invertible matrix $M$ by a random sparse matrix $A$, we have that the corresponding perturbation of the inverse $\left(M+A\right)^{-1}-M^{-1}$ is typically less sparse than $A$.
This claim can be used to show that in natural settings, if we want to condition e.g. on $f^{\rightarrow}$ being sparse, this will result in an $f^{\leftarrow}$ that is comparatively less sparse, and vice versa. In 3.2.2 below, we propose a communication setup where the sparsity of $f^{\rightarrow}$ is naturally favored, yielding a FW AoT as observed in the natural languages (see 2.2 above).
3.2.2 A Simple Communication Setup
In the previous section, we have shown that the emergence of an AoT is natural in the sparse setting: if we e.g. condition $f^{\rightarrow}$ to be sparse, this will yield an inverse $f^{\leftarrow}$ that is less sparse. To motivate the importance of sparsity, and in particular of FW sparsity (for the presence of a FW AoT), we give a simple communication setup.
Suppose Alice and Bob are (human) agents with FW predictors having learned a common language $\mathbb{P}_{B}$, and Carol is an (alien) agent with a BW predictor having learned $\mathbb{P}_{B}$ as well. Now suppose Alice wants to teach Bob a new language $\mathbb{P}_{A}$ by sending him samples from $\mathbb{P}_{A}$; how easy this is will typically depend on how far away $\mathbb{P}_{A}$ is from $\mathbb{P}_{B}$, i.e. how sparse $f_{BA}^{\rightarrow}=f_{A}^{\rightarrow}-f_{B}^{\rightarrow}$ is. Assume Alice is only able to teach $\mathbb{P}_{A}$ to Bob if $f_{BA}^{\rightarrow}$ is sparse enough (note that Alice needs to learn $\mathbb{P}_{A}$ herself, it is reasonable to assume that she will only be able to do so if $f_{BA}^{\rightarrow}$ is sparse enough). Conditioning on Alice being able to teach $\mathbb{P}_{A}$ will hence yield (with high probability) a FW AoT: following Claim 8 above, $f_{BA}^{\leftarrow}=f_{A}^{\leftarrow}-f_{B}^{\leftarrow}$ will be typically less sparse than $f_{BA}^{\rightarrow}$ and $\mathbb{P}_{A}$ will be harder for Carol to learn than for Bob. This will ultimately impact the language structure: if e.g. $f_{BA}^{\leftarrow}$ was often sparser than $f_{BA}^{\rightarrow}$, it would be profitable to ‘restructure’ the language, expressing $y\leftrightarrow x$ rather than $x\leftrightarrow y$. This suggests that selection pressure may cause languages to evolve to take a form where $f_{BA}^{\rightarrow}$ is often sparser than $f_{BA}^{\leftarrow}$ , yielding a consistent FW AoT.
3.2.3 Experimental Results
In this section, we present experimental results supporting the claims of the previous sections. We first consider a dataset made of linear languages with $x,y\in\mathbb{F}_{2}^{25}$, for different sparsities of $f^{\rightarrow}$. We train a GPT-1 model on these datasets (see Appendix B.2) and plot the final losses in Fig. 3, confirming that sparser matrices are easier to learn.
Figure 3: Models loss at the end of training vs $f^{\rightarrow}$ sparsity.
In the second experiment, we consider a model’s ability to learn a sparse update, given a learned prior: we first train FW/BW models on a linear language with $m=20$ (with a sparse FW matrix) until it is learned perfectly, then generate a new linear language by a sparse perturbation of the learned language. Table 5 shows the losses of both models after 400 gradient descent steps. Again, we see that the FW model adapts better to the sparse modifications (see Appendix B.2.3).
Table 5: GPT-1 losses for different perturbations of the learned prior.
2-bit flips 4-bit flips 6-bit flips
FW $0.347\pm 0.009$ $0.354\pm 0.007$ $0.367\pm 0.014$
BW $0.353\pm 0.008$ $0.371\pm 0.011$ $0.387\pm 0.013$
3.2.4 Non-Linear Case and Extensions
If we consider a more general model of languages compared to Section 3.2.1, based on arbitrary functions $f:\mathbb{F}_{2}^{m}\to\mathbb{F}_{2}^{m}$, the notion of sparse map needs to be adapted to consider sparse circuits, made of relatively few logic gates (AND/OR/XOR) in the linear circuits. We can expect that for a random sparse circuit model $f^{\rightarrow}$, the inverse (or any pre-image computation) will typically be much less sparse. This is suggested by the fact that inverting circuits is expected to be computationally hard (roughly the content of the $P\neq NP$ conjecture). We can hence expect an AoT for similar reasons. Due to the computational hardness of inverting nonlinear circuits, a difference with the linear case can be expected to arise: the nature of the AoT in this case can also be of Type 1 (see Section 3); in some cases, a BW model may simply be unable to represent what the FW model learns, as in the example of Section 3.1.
4 Discussion
In this paper, we have investigated the abilities of auto-regressive LLMs, which predict tokens sequentially: for a given measure (dataset), we compare the abilities of two models (FW/BW). Theoretically, if both models learn to represent the same measure, their average log-perplexities coincide. Somehow surprisingly, we discover the existence of an Arrow of Time (AoT) for natural language datasets: across a wide variety of models and hyperparameters, we see that all else being equal, FW models consistently exhibit a lower perplexity than BW ones; we see this difference emerge as soon as the model is large enough; its causes appear to lie in long-range correlations in the data, as the magnitude of this effect increases with the context length. We then propose a framework to explain this phenomenon, based on computational complexity and sparsity ideas: we construct examples of synthetic datasets based on arithmetic operations that display an asymmetry in terms of computability (despite being information-theoretically reversible); finally, we propose a setup where a FW AoT like the one seen in natural language can spontaneously emerge. Our work suggests a number of possible future research directions:
  • Are AoTs universal across all human languages?
  • Are there AoTs in other types of languages, e.g. computer code, binaries, DNA code, or bitmap files?
  • How to explain the variation in magnitude of the AoT across languages?
  • Are there AoT scaling laws with respect to model sizes?
  • Are natural language AoTs of Type 1 or 2 (in the sense of Section 3)?
  • For very long training times, is there a difference between train and test AoTs?
  • Can AoTs be revealed in continuous settings, e.g. for video prediction?
  • Is there a link with other AoTs, e.g. in thermodynamics?
  • Can we generalize the idea of flipping the order of the tokens to other permutations of the context window?
  • Are AoT and computational hardness deeply linked?
In conclusion, the concept of AoT appears to be related to subtle properties of natural language data, revealed through their interplay with autoregressive LLMs. This idea’s applicability seems wide, and a promising new tool to reveal the presence of deep structural features in data. Further study of its theoretical origin could prove fruitful towards uncovering links between AoTs and complexity theory.
Acknowledgements
The authors would like to thank Stéphane d’Ascoli, Gloria Capano, Diego Dorn, Franck Gabriel, Ron Maimon, Jacob Menick, Christos Papadimitriou, João Penedones, Matthieu Wyart, Nicolas Zlatoff for interesting discussions. Support from the Blavatnik Family Foundation, the Latsis Foundation, the NCCR SwissMAP, and from an EPFL FSB Seed Funding Grant are gratefully acknowledged.
Appendix A Details on training on natural languages
In this appendix, we provide more details on the training of our models on natural languages.
For any experiment, the precise hyperparameters that were used can be found in the supplementary material, under the folder “Training Parameters”, in .$json$ format. Those files can also be used to reproduce any experiment using the codebase, as explained in the README.md of the Github repo.
All experiments (save for the 512 context size) were run on a single A100 GPU, adjusting the batch size to fit in memory.
Concerning the shuffling of the dataset, we proceed as follows for all the experiments. We begin by separating the textual dataset in ‘sentences’ of the appropriate context size $n$, with a stride of $n/2$ (i.e., if we have a context size of 4 and the text is ABCDEF, this results in two sentences, ABCD and CDEF). This is to ensure that all tokens appear in the training data with at least some context. After that, we shuffle the obtained sentences with a set seed. We withhold 250k sentences (1000 batches at 250 batch size) for validation. The inversion of the tokens is made at the level of each batch; in this way, when training, the FW/BW models see the data in the same order, preventing the emergence of undesirable differences.
A.1 Model sizes
Table 6 provides more detail on the model sizes used in the paper.
Table 6: Model sizes used throughout the experiments. $d_{embed}$: number of embedding dimensions. $n_{heads}$: number of attention heads. $n_{layers}$: number of transformer blocks (attention + MLP). parameters : total number of parameters, including the last linear layer which projects on vocabulary size (commonly referred to as the ‘head’).
GPT2 model name $\rightarrow$ Nano Micro Mini Small GPT1 Medium XL
$d_{embed}$ 48 128 192 380 768 1024 1600
$n_{heads}$ 3 4 6 10 12 16 25
$n_{layers}$ 3 4 6 10 12 24 48
parameters 4.92M 13.7M 22.0M 55.6M 162M 405M 1.6B
In the MLP layer, all models have one hidden layer with $4*d_{embed}$ hidden dimensions, a.k.a. an MLP ratio of $4$.
A.2 Different languages
A.3 Context Window size
For testing for the influence of the context window length, we use a GPT2-Medium model, with a context window length going from $16$ to $512$ tokens. Because of the different context lengths, it will take a model with a small context length many more gradient steps to see the same amount of data. For this reason, we do not train all models up to the equivalent of 1 epoch of the French dataset ($\sim 26.6B$ tokens), but rather train them for sufficiently long so that perplexity differences stabilize, and that their losses converge. Due to the cosine annealing learning rate schedule, we stop the training at the end of a cosine decay, to avoid the bump in the loss caused by a warm restart (note however that this has little to no effect on the perplexity differences).
In Table 7, we record the number of steps (i.e. minibatches) that were seen for each context length.
Table 7: Number of batches seen during training for the different context lengths. Not that the recorded batch size is “effective”, potentially obtained through aggregation of smaller batch sizes.
Context length $\rightarrow$ 16 32 64 128 256 512
batch size 500 500 500 400 180 156
seen batches 1.74M 1.62M 1.64M 1.02M 0.5M 0.6M
For this experiment, because of our memory limitations, we trimmed down the English dataset, which was too big when working with context windows of lengths $16$ and $32$. Note that the 256 context length experiment in this section is thus slightly different from the one recorded in Table 3, due to the different datasets (as well as batch size and cosine annealing period).
A.4 Other Models
The GRU and LSTM implementations we used for these experiments are those natively implemented in the pytorch.nn of the Pytorch Python package. In each case, we choose the “input size” (i.e., the number of embedding dimensions for each token) to be equal to the “hidden size” (i.e., the number of hidden dimensions in the RNN’s hidden state).
Table 8: Parameters for different sizes of LSTM and GRU models.
Small Medium Large
$hidden_{dim}$ 256 512 768
$n_{layers}$ 1 3 5
Although RNNs have technically no limit on the context length, for training purposes, to allow for a backward pass, we feed them with batches of texts of lengths $256$.
A.5 Other Languages
For the experiments on other languages, we decided to stop the training at the equivalent of 1 epoch of Greek, which was one of the smallest datasets on which we were training (along with Turkish). This choice was made so for all languages, models never saw the same datapoint more than once (note that we trained on Greek for two epochs, as shown in Fig. 5; this suggests that our results remain valid beyond one epoch).
For GPT2-Medium, a batch size of $90$ was used, with one warm restart during training. For the XL models, a batch size of $26$ was used, aggregated $6$ times for an effective batch size of $156$, which didn’t allow for a warm restart before the one epoch of Greek.
In Fig. 4 we showcase the validation loss of all the language during training, displaying the same behaviour as observed for English and French in section (2.2.1).
Figure 4: Validation losses through training of GPT2-Medium for different languages, both in FW and BW. Note that we clipped the beginning of training to better showcase the subtle differences in loss.
For completeness, we also attempted to run the training for Greek for 2 epochs, to see if memorization of the dataset may affect the arrow of time. In Fig. 5, we display the validation loss during training. Comparing the difference in performance at 1 and 2 epochs, it remains almost exactly the same. It would be interesting to test this with bigger models, and several epochs of training.
Figure 5: Validation loss for two epochs of training on the greek dataset, for forward and backward models.
A.6 BPE Tokenization
For the tokenization, we use the Huggingface implementation of the BPE Tokenizer (link), using the method Tokenizer.train_from_iterator. The tokenizers are trained on the same CC100 dataset on which we train the model.
For the experiment in which we train the BPE tokenizer in reverse (see Section 2.2.6), we first reverse the Greek dataset at the character level (not at the byte level, as this would make the output of the model unreadable), then train the BPE tokenizer on this new dataset. We then train a FW and a BW model on this character-flipped dataset, tokenized with the new BPE tokenizer. We will call a model ‘backward’ (BW) if it processes tokens in the opposite order w.r.t. the natural reading direction (hence the ‘previous-token predictor on the character-flipped dataset’ corresponds to what we will call the FW model). Thus, if the arrow of time is a property of the language and not a tokenization artifact (as we expect), we expect both BW and FW models to perform similarly, independently of the tokenization scheme. Fig. 6 confirms this; relying on the reverse BPE tokenization introduces almost no difference, except for a very small discrepancy for the BW models. Since it produces no noticeable difference for the FW model, this small difference in the BW case is likely due to initialization.
Figure 6: Validation curves for models training on the Greek dataset on forward and backward (label prefixed with ‘rev’) BPE tokenizations. The arrow of time remains in the same, forward direction, despite the different tokenization schemes.
Appendix B Linear Language Toy Model
B.1 Matrix Inverse Sparsity
Here, we substantiate the assertions of Claim 8 by running numerical experiments. To this end, we wish to look at $n\times n$ matrices in $\mathbb{F}_{2}$ with a given number of non-zero elements (which we will call $k$), and compute $k$ in the inverse. To generate invertible matrices with extremely high sparsity (low $k$), we proceed as follows. We start with the identity matrix $Id$, which is the only invertible matrix (up to permutations, which do not affect sparsity) when $k=n$. To generate a matrix with approximately $k$ non-zero elements, we flip $k-n$ elements of $Id$ at random. This allows us to often obtain invertible matrices when $k$ is close to $n$, which would not be the case if we simply selected the $k$ non-zero elements at random. When $k$ becomes bigger than $n$, the initial presence of the identity is quickly erased.
We choose $n=30$, generate matrices with $k=[30,250]$, and record the average number of non-zero elements in the inverse, given $k$ fixed. Fig. 7 clearly confirms that the sparsity of the inverse is generally much lower.
Figure 7: Plot displaying the connection between the sparsity of a matrix and its inverse. It is clear that on average, the inverse of a sparse matrix is less sparse. The number of non-zero elements for the inverse caps at 450450~{}450450
B.2 Linear Languages Experiments
In this section, we give more details on the experiments of Section 3.2.3.
B.2.1 Linear Language Dataset
Given a matrix $M$ of size $n\times n$, the associated linear language dataset will contain sentences of $72n+$ 7 tokens in the form $x\_\_\_\_\_\_\_y$, where $x,y\in\mathbb{F}_{2}^{n}$, and the underscores are added as padding, providing the model with more tokens if needed to perform more complex computations. The vector $x$ is drawn at random, and $y$ is computed with $y=Mx$. We finish by randomly flipping each bit with probability $p=0.01$ (which we call ‘adding’ perturbations), aimed at smoothing out the probabilities output by the model. This is necessary for the ‘fine-tuning’ experiments, as otherwise the models become too confident in their predictions, and any small change $M$ results in a huge change in the loss, leading to a catastrophic forgetting of the learned prior.
B.2.2 Sparsity Levels
In the first experiment, we generate a linear language model in $\mathbb{F}^{25}_{2}$, with $p=0$ perturbations, and matrices with a number $25+k$ of non-zero elements, where $k\in[0,2,4,8,10,14,18,20,25,30,35,40,45,50]$. We then train a transformer model of size GPT1 (see Table 6) on $600k$ sentences, with batch size $200$. Note that the context size of the model matches exactly the number of tokens in one sentence. Final losses are reported in 3. Note that for lower sparsities, the trend is not obvious: this is due to the high variance in the final learning rate, as the learning of only a few non-zero elements is binary, depending on the initialization, the model either learns the matrix perfectly quickly, or it usually struggles to find the last few non-zero elements. Fig. 8 displays this behavior in the case $k=4$. Note that the perturbations somewhat reduce this behavior, but don’t cancel it completely.
Figure 8: Average perplexities for each token in the linear language after $600k$ sentences, for two runs with $k=4$. One of the models is very close to the optimal solution, while the other is missing a single token. It usually takes a long time for the model to correct this, leading to higher variance in the final losses.
Fig. 9 displays typical learning dynamics for this problem, for $k=8$ (high sparsity) and $k=40$ (medium sparsity). We remind that in principle, given a large enough model, and enough training steps, the model should be able to find the optimal solution (hardness of type (2), see Section 3).
Figure 9: Loss during learning for $k=8$ and $k=40$ sparsities. The first plateau simply arises when the model learns to guess all coefficients randomly. Subsequently, the $k=8$ experiences plateaus each time it learns more non-zero coefficients. The learning of $k=40$ is much smoother, as discovering non-zero coefficients doesn’t lead to perfect predictions right away.
B.2.3 Sparse updates
In the second experiment, we choose a linear language in $\mathbb{F}^{20}_{2}$. We generate the dataset in the same way explained in Section B.2.1. We begin by training a FW model and a BW one on this language, until both models learn it almost perfectly (batch size 200). For this reason, we choose a very sparse matrix, with $k=6$ (in this specific example, the inverse has $k=10$). As expected, this takes much longer for the BW model, as displayed in Fig. 10.
Figure 10: Training loss for the FW and BW models, trying to learn a linear language with a matrix $k=6$. While the FW model learns the quasi-optimal solution very quickly, the BW model remains stuck on a plateau for a long time. This corresponded to a single element of the predicted vector which was missing, as in Fig. 8.
Once this ‘prior’ is learned, we generate perturbation of the learned matrix by flipping $e$ entries of the matrix randomly, conditioned to the fact that it should remain invertible. We then train the models further on this new dataset, for a relatively small amount of steps (400 gradient steps). We also lower the learning rate to $8\times 10^{-6}$, with 10 steps of warmup, again to prevent catastrophic forgetting of the prior. The training dynamics are displayed in Fig. 11, in the case $e=4$.
Figure 11: Averaged loss for forward and backward models, when trying to learn a sparse forward perturbation of the Linear language. In the first $\sim 100$ steps, the curves are similar as they simply reduce their confidence on the new, perturbed tokens to random chance. Subsequently, they begin learning the perturbation. The forward model is clearly at an advantage.
We observe that the FW model adapts better than the BW one, and this is due to the fact mentioned in Claim 8, namely that a sparse FW update will generically result in a less sparse BW update.